OpenStudy (anonymous):

solve the equation 2[(sin^2 x)(cosx)]=cosx by factoring. give the solutions on [0, 2pi]

OpenStudy (anonymous):

\[2(\sin ^{2}x)(cosx)=cosx\]

OpenStudy (jdoe0001):

\(\bf 2[sin^2(x)cos(x)] = cos(x) \implies sin^2(x) = \cfrac{\cancel{cos(x)}}{2\cancel{cos(x)}}\\ sin(x) = \sqrt{\cfrac{1}{2}} \implies sin(x) = \cfrac{1}{\sqrt{2}} \implies sin(x) =\cfrac{\sqrt{2}}{2}\)

OpenStudy (jdoe0001):

\(\bf sin(x) = \pm \cfrac{\sqrt{2}}{2} \implies sin^{-1}(sin(x)) = sin^{-1} \left(\pm \cfrac{\sqrt{2}}{2}\right)\)

zepdrix (zepdrix):

Subtracting cosine from each side,\[\Large 2\sin^2x \cos x-\cos x=0\]Factoring a cosine out of each term,\[\Large \cos x(2\sin^2x-1)=0\] Doe's method works also, but we end up losing a solution because of the cancellation of those cosines :c

OpenStudy (jdoe0001):

hmm, I see that

OpenStudy (anonymous):

ok, so with either way how do you find the other solutions, or is there only one? i forget how the renaming the angles works on the unit circle

zepdrix (zepdrix):

renaming the angles? :o

OpenStudy (anonymous):

i dont know what its called but isnt there something about having to change them to fit in the right quadrant or something im not really sure

zepdrix (zepdrix):

\[\Large \cos x(2\sin^2x-1)=0\]Using this method, we would set each factor equal to zero and solve for the trig functions separately.\[\Large \cos x=0 \qquad\qquad\qquad 2\sin^2x-1=0\] Where is cosx=0? Recall that cosine is the x component of our angle. The x-coordinate is 0 when we're up at pi/2 and down at 3pi/2. So those are 2 of our solutions. \(\large x=\pi/2\) and \(\large x=3\pi/2\)

zepdrix (zepdrix):

right quadrant? Hmm we might have to worry about that when we deal with the sine factor, let's see...

zepdrix (zepdrix):

\[\Large 2\sin^2x-1=0\]Solving for sin^2x gives us,\[\Large \sin^2x=\frac{1}{2}\]And then solving for sinx, taking the root of both sides gives us,\[\Large \sin x=\pm\frac{\sqrt{2}}{2}\]Hmm

OpenStudy (anonymous):

ok so the \[+\frac{ \sqrt{2} }{ 2 } \] would work, but the negative one wouldnt because it isnt within 2pi so would there only be those three answers?

OpenStudy (anonymous):

or would pi/4 just be the answer for the second part

zepdrix (zepdrix):

The negatives ARE in 2pi, right? :) I was just trying to think about whether or not they would actually work for a solution. Since the sinx is being squared, it looks like all 4 angles will work. \[\large x=\pi/4, \qquad x=3\pi/4, \qquad x=5\pi/4\qquad x=7\pi/4\] Lemme test one just to make sure I'm not being silly.

zepdrix (zepdrix):

\[\large x=5\pi/4 \qquad\to\qquad 2\sin^2(5\pi/4)\cos(5\pi/4)=\cos(5\pi/4)\] Soooo this gives us,\[\large 2\left(-\frac{\sqrt2}{2}\right)^2\left(-\frac{\sqrt2}{2}\right)=\left(-\frac{\sqrt2}{2}\right)\]Which simplifies to,\[\large 2\left(\frac{1}{2}\right)\left(-\frac{\sqrt2}{2}\right)=\left(-\frac{\sqrt2}{2}\right)\] Yah that seems to work ok :)

zepdrix (zepdrix):

A little rusty on your trig angles maybe? :D

OpenStudy (anonymous):

haha yeah :)

zepdrix (zepdrix):

Bahhh not the best drawing.. i was trying to show that the sine represents the vertical (y component) of that angle. So you can clearly see when the y value is negative for those 2 lower angles.

zepdrix (zepdrix):

Ehh maybe that's more confusing XD lol. Just brush up on your trig!! :3

OpenStudy (anonymous):

so there are six answers to this question?

zepdrix (zepdrix):

Here maybe this will help :) https://www.desmos.com/calculator/fq8u4tetb1

zepdrix (zepdrix):

Solutions means ~ points of intersection. How many times do they intersect between 0 and 2pi? I tried to label the x-axis in radians, so it'd be easier to read. Hopefully that feature is showing up correctly.

OpenStudy (anonymous):

ok thank you so much for all of your help with this!!

zepdrix (zepdrix):

np c: I find trig to be a little bit difficult to explain without going into a ton of detail. So maybe just look back at a unit circle or something. I hope you can make sense of those 6 solutions :[